# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ IV-E Graph-Based Mobile Vision
Computer vision tasks accept inputs in the forms of images, videos, and point clouds, where many of them are representable by graphs. For instance, Fig. 7(g) shows an image of the digit “6” in 18$\times$12 pixels, where by viewing pixels as vertices and their direct adjacency as links, it can be recognized as a graph. Another example is the skeleton-based graphs in many images of human actions, as discussed in Sec. IV-C2 and depicted in Fig. 7(e). GI models are applied to these graphs to assist various traditional computer vision tasks like image classification, object detection, point cloud analysis, etc.
IV-E1 Image Processing
@@
In the evolving landscape of image processing, Graph Inference (GI) models have emerged as a powerful paradigm, particularly for tasks that require an understanding of the relationships and context within visual data. At the forefront of this advancement is the work by Liu et al. [174], who introduced the Structure Inference Network (SIN). This pioneering approach in object detection harnesses the power of graph structures to model the scene context and object relationships. By treating objects as nodes and their interactions as edges within a graph, SIN elevates the traditional object detection algorithms, providing a more holistic and contextual understanding of the visual scene. Expanding on this foundation, Chen et al. [175] delve into the domain of multi-label image recognition. They propose a GCN model that constructs a directed graph of object labels, capturing the intricate dependencies and co-occurrences among them. This model not only builds upon the graph-based reasoning initiated by Liu et al. but also extends the concept to the challenges of recognizing multiple objects within an image. Following this thread of innovation, Lu et al. [176] first apply the GCN to image semantic segmentation with their Graph-FCN, an approach that redefines the task as a graph node classification problem, successfully integrating local location information often overlooked by standard deep learning methods.
IV-E2 Video Analytics
@@
Video analytics with GI models are mainly for HAR and Multi-Object Tracking (MOT). While HAR has been thoroughly explored in Section IV-C2, this section delves into the latter, focusing on the innovative approaches that have emerged. Braso et al. [177] set the stage by introducing a groundbreaking graph-based framework that employs message-passing networks (MPNs). Their approach not only integrates learning within the MOT paradigm but also crucially extends it to the data association phase, which has traditionally been a bottleneck. Simultaneously, Liu et al. [178] introduce the graph similarity model for MOT, which integrates a unique graph representation of object features and relationships, coupled with a graph matching module, to enhance robustness against occlusion and similar appearances in tracking scenarios. Both Brasó et al. and Liu et al. have independently demonstrated the potential of GI models to transform MOT by focusing on different yet complementary aspects of the problem—global reasoning and robust feature representation. Building upon these foundational works, Zhang et al. [179] later introduced a method that combines spatio-temporal correlation learning with GNNs. This approach dynamically constructs group structures and enhances node feature aggregation, effectively addressing the challenges of MOT in terms of accuracy and efficiency.
IV-E3 Point Cloud Processing
@@
Different from images in 2D representations, point clouds are organized in 3D structure and provide richer semantics in describing scenes and objects. For instance, many autonomous vehicles use LiDAR to scan the environment and export point cloud data for perception, as in Fig. 8. Yet the 3D data are still accommodated in a graph representation, and thereby GI models are leveraged for efficient processing. Wang et al. [180] introduce EdgeConv, which enhances the representation power of point clouds by recovering topological information. This method makes use of dynamic graph construction to capture local geometric features, proving its suitability for high-level tasks in computer graphics and vision. Building on this idea of dynamic feature extraction, Shi et al. [173] propose Point-GNN, a GI model for object detection from LiDAR point cloud, which incorporates an auto-registration mechanism to reduce translation variance and a box merging and scoring operator for accurate detection over multiple vertices. Extending these advancements, Lin et al. [181] develop 3D-GCN to address the challenges of processing and summarizing information in unstructured point clouds for 3D vision applications. Their approach is particularly effective when dealing with data variations such as shift and scale changes, aligning with the principles seen in EdgeConv. Furthermore, Zhang et al. [182] present the linked dynamic GNN, which tackles the challenge of processing sparse, unstructured point clouds by linking hierarchical features from dynamic graphs. This method optimizes the network’s architecture for more effective point cloud classification and segmentation, showing how linking hierarchical features can avoid the vanishing gradient problem and improve overall performance.
V Edge Computation for Graph Intelligence
While graph intelligence enables edge networks, edge networks serve as the infrastructural support to GI applications and optimizations, i.e., allowing edge computation of GI models. Compared to traditional cloud execution of GI models that fully sends data to the cloud datacenter for centralized processing, edge model computation reserves a majority of data at the network edge, rendering localized, low-latency, and privacy-preserved services around end users. Nonetheless, given the multi-tier hierarchy of edge networks, edge computation of GI models can be flexibly provisioned across edge platforms, e.g., through edge-cloud federated execution, on-device computation, and distributed edge collaboration.
V-A Federated Graph Learning
@@
Edge-cloud synergy extends cloud resources as remote computing assists in optimizing the performance and efficiency of user-centric computing and data processing. Specifically, it enables certain computing tasks reserved at the edge for immediate processing, while employing the cloud for offloading more resource-intensive tasks with long-term storage or/and complex data analysis. This hybrid computing paradigm combines the differentiated capabilities of edge and cloud computing, thereby admitting flexible service computing for various applications [183]. For edge computation of GI models, typical workloads in the form of edge-cloud synergy are FL of GI models, i.e., Federated Graph Learning (FGL). Note that FGL is distinct from GI-assisted FL (GFL) as discussed in Sec. VI-D. Although they both combine GL and FL, the targeted model to be federatively trained in FGL is exactly GI models but GFL allows the targeted model to be any model not only GI models. Readers particularly interested in the combination of GI and FL may find dedicated reviews [184, 185] for more information. With respect to the segmentation of the data space, FGL can be divided into horizontal ones (graph topology space) and vertical ones (feature data space).
@@
V-A1 Horizontal Federated Graph Learning
@@
The first setting considers FGL with explicit links across subgraphs on FL clients, as in Fig. 9. The conventional approach in this setting entails training local GI models within individual clients to initially acquire the local representations or node embeddings of the respective graphs. Subsequently, a Federated Learning (FL) algorithm is applied to facilitate the aggregation process and the FL server, which is usually a cloud server, accumulates the model parameters or gradients received from the clients and performs FL aggregation. The resultant updated parameters are then returned to the clients for subsequent rounds of training. The introduction of FGL within this framework is contingent upon the specific research problems they address. For Non-IID graph data issues, many approaches intend to enhance the adaptation performance of local models or acquire a robust global model through model adjustment like model interpolation [186, 187], loss regularization [188], and meta learning [189]. Data adjustment approaches are also frequently raised to address that, which aims at eliminating data heterogeneity using methods like sample reweighting [190] and manifold learning [191].
The second setting assumes that certain links connecting vertices across different clients are implicit and absent, illustrated in Fig. 9. For the case where the implicit links connect vertices with dissimilar IDs, a prevalent strategy involves rectifying the local graphs by reconstructing the missing links. This ensures the local graph’s completeness, thereby facilitating high-quality graph representation. Furthermore, the inclusion of links connecting clients aids in mitigating the challenges associated with non-IID data. FASTGNN [192] introduces a straightforward link generator in the FL server, which is responsible for reconstructing absent links between clients by introducing Gaussian randomly generated links. These reconstructed edges are broadcast to all clients during FL training, prompting them to update their local graphs. In a similar vein, the link generator in FedGL [58] produces a global pseudo graph with vertex embeddings provided from clients. Subsequently, this global pseudo-graph is disseminated to all clients, enabling them to amend their local graphs to facilitate GI model training. Based on the simple yet effective idea of link generation, follow-up works further develop more complex techniques for better performance [193, 194], consider privacy issues with security means [195, 196, 197], and explore to improve communication efficiency [198, 199].
@@
V-A2 Vertical Federated Graph Learning
@@
In the initial scenario, graph topology, vertex features, and vertex labels are distributed across different clients. As a result, clients do not have access to the complete graph data, which includes both vertex features and graph topology, as illustrated in Fig. 10(a).
@@
In the first case, graph topology, vertex features, and vertex labels are distributed across different clients. Consequently, clients may not be able to obtain the full graph data with respect to both vertex features and graph topology, as shown in Fig. 10(a). To orchestrate these clients in such semi-blind situations, SGCNN [205] computes a similarity matrix based on the dynamic time warping algorithm to sketch the graph’s topology without knowing the underlying structure. One-hot encoding is employed to describe vertices’ features under privacy requirements, which will be transmitted to the server to train the global GI model for GI-based tasks. FedSGC [206] considers a two-client decentralized setting with vertices’ features and graph topology being respectively possessed. In this scenario, FedSGC proposes an additively Homomorphic Encryption (HE) method to ensure privacy preservation during FGL.
@@
The second case entails the ownership of solely disparate vertex feature spaces by clients, while the graph topology remains accessible to all clients (Fig. 10(b)). For the case with only two clients and a central server, FedVGCN employs HE to encode intermediate data transferred across clients, where the server generates encryption key pairs and performs global FL aggregation. FMLST [207] assumes a shared global spatio-temporal pattern across clients and applies a MLP to fuse local pattern and global patterns for personalizing client models. Graph-Fraudster [208] investigates adversarial attacks on this vertical FGL scenario and shows differential privacy mechanisms and top-k mechanisms as two viable defense strategies against such attacks.
V-B On-Device Computation
V-B1 Resource-Efficient Procedure
Without modifying GI models, resource-efficient procedure design indicates the endeavors that develop GI processing flows in minimal resource usage, e.g., memory footprint. For instance, centering on the out-of-memory issues where GI models’ sizes are too large to fit into edge devices’ memory, Zhou et al. [209] design a feature decomposition approach that decomposes the dimension of feature vectors and performs aggregation respectively, reducing up to 5$\times$ memory footprint against generic GI frameworks. Building on the theme of memory efficiency, SUGAR [210] targets resource-efficient training on resource-constrained edge devices through a subgraph-level training scheme that first partitions the initial graph into a set of disjoint subgraphs and then performs local training for individual subgraphs. Complementing these approaches, RAIN [211] explores the opportunity of conservatively processing similar inference batches based on local sensitive hash and reusing repeated vertices’ data among successive batches to reduce redundant data loading.
There are also some works that utilize channel pruning methods to reduce input feature dimensions for GI model inference acceleration. Zhou et al. [212] set the stage by employing a LASSO regression-based pruning technique, which adeptly identifies the most influential feature channels, thus ensuring that the pruning process is both effective and minimally disruptive to the model’s accuracy. Yik et al. [213] extend this line of inquiry by conducting a meticulous analysis of the trade-offs inherent in channel pruning. They develop algorithms that not only preserve the model’s predictive prowess but also enhance computational efficiency, aligning seamlessly with Zhou’s foundational work by further honing in on the most valuable features. In a complementary contribution, the STEP framework [214] addresses the unique challenges posed by dynamic graphs. By proposing an unsupervised approach to graph pruning, it builds upon the principles established by its predecessors, focusing on the timely removal of redundant links. This ensures that the graph’s structure remains both informative and agile, catering to the needs of evolving datasets.
V-B2 Resource-Efficient Model
To automate the resource-efficient model design for edge platforms, some researchers [218, 219, 220] introduce NAS techniques to the GI domain and present hardware-aware NAS frameworks tailored to GNNs, aiming to strike a balance between model accuracy and efficiency corresponding to the characteristics of targeted devices. Zhou et al. [218] present HGNAS, a hardware-aware graph neural architecture search (GNAS) framework that optimizes GNN designs for edge devices by balancing accuracy and efficiency using a hardware performance predictor. Similarly, PaSca [219] constructs scalable GNNs through a novel scalable GNAS paradigm, enabling systematic exploration of a vast design space to optimize GNNs for both accuracy and efficiency via multi-objective optimization. Building upon these concepts, MaGNAS [220] is a mapping-aware GNAS framework that efficiently processes vision-based GNN workloads on heterogeneous multi-processor platforms by identifying optimal GNN architectures and mappings for maximized resource efficiency and performance trade-offs. Together, these works highlight the growing trend towards hardware-aware and scalable GNAS frameworks that address both the accuracy and efficiency requirements crucial for deploying GNNs on edge devices.
V-B3 Model Compression
Compressing a complete model into a smaller one that appropriately fits the capability of the targeted edge device has been a popular deployment means for many edge intelligence services. The smaller model expected, which is usually with fewer parameters, allows much less computing cost and thus significant acceleration for on-device computation. To obtain such a lightweight model for EGI, researchers typically employ three ways, i.e., model quantization, knowledge distillation, and neural architecture search.
Model quantization intends to quantize GI model parameters from an origin, larger bitwidth to a targeted, smaller bitwidth and thus reduce the total size of the model. In this context, SGQuant [221] proposes a multi-granularity method that can quantize feature vectors in the component level, topology level, and layer level to meet diverse data precision demands. It also designs an automatic bitwidth-selecting algorithm to attain proper configuration among different quantization granularities. Building upon the quantization paradigm, Degree-Quant [222] pioneers a quantization-aware training scheme specifically crafted for GI models. This scheme is designed to facilitate low-precision inference on devices with constrained resources, without compromising the model’s predictive capabilities. VQ-GNN [223] takes a quantum leap in this domain by introducing a unified framework that applies vector quantization (VQ) for dimensionality reduction and learns a small number of quantized reference vectors of global node representations, which viably avoids the “neighbor explosion” problem and enables faster mini-batch training and inference of GI models. Following this quantization principle, some researchers go one step further with GI model binarization, i.e., binarizing parameters in GI models for aggressive execution speedup [224].
Knowledge Distillation (KD) aims at extracting knowledge from a complex teacher model and transferring it into a compact student model while maintaining comparable performance [225]. While KD is originally applied in the computer vision domain [226, 227, 228], Yang et al. [229] first introduce KD for GI models along with a local structure preserving method for teacher-student similarity measurement and experimentally show its effectiveness in GI-based 3D object detection tasks. Along with this line, Yang et al. [230] combines parameterized label propagation and feature-based MLP in the KD’s student model, allowing better utilization of the knowledge from both the teacher model and prior basement. TinyGNN [231] focuses the KD principle on accelerating model inference and designs a peer-aware module and a neighbor distillation strategy for preserving model accuracy.
V-C Distributed Edge Collaboration
@@
V-C1 Edge-Cloud Collaboration
@@
Edge-cloud collaboration differs from traditional integrated cloud offloading by splitting computation workload between edge platforms and cloud servers and thus can fortify the duty of the edge side and alleviate the reliance on the remote cloud [17, 236]. This allows edge facilities to be a key processor that blocks sensitive raw data from the cloud and eliminates redundant data transmission in the backhaul. Following this technical path, Branchy-GNN [237] concentrates on GI-based point cloud processing and applies edge-cloud collaboration with a learning-based source-channel coding scheme for bandwidth-efficient intermediate data transfer. It also borrows the idea of joint split learning and early exiting to reduce the computational cost on edge devices [17, 238]. Pyramid proposes a hierarchical architecture for edge-cloud synergetic GNNs that trains a local, small model on edge servers with its collected data and a global model on a cloud server to capture inter-region spatio-temporal relationships. During the runtime, Pyramid [239] allows GI-based predictions with either the local model or the global model in an on-demand fashion and achieves lower inference latency against baselines.
V-C2 Multi-Edge Collaboration

### v2

@@ V-E Graph-Based Mobile Vision
Computer vision tasks accept inputs in the forms of images, videos, and point clouds, where many of them are represented by graphs. For instance, Fig. 7(g) shows an image of the digit “6” in 18$\times$12 pixels, whereby viewing pixels as vertices and their direct adjacency as links, it can be recognized as a graph. Another example is the skeleton-based graphs in many images of human actions, as discussed in Sec. V-C2 and depicted in Fig. 7(e). GI models are applied to these graphs to assist various traditional computer vision tasks like image classification, object detection, point cloud analysis, etc.
V-E1 Image Processing
In the evolving landscape of image processing, Graph Inference (GI) models have emerged as a powerful paradigm, particularly for tasks that require an understanding of the relationships and context within visual data.
@@ V-E1 Image Processing
At the forefront of this advancement is the work by Liu et al. [163], who introduced the Structure Inference Network (SIN). This pioneering approach in object detection treats objects as nodes and their interactions as links within a graph, providing a more holistic and contextual understanding of the visual scene. Expanding on this foundation, Chen et al. [164] delve into the domain of multi-label image recognition. They propose a GCN model that constructs a directed graph of object labels, capturing the intricate dependencies and co-occurrences among them. This model not only builds upon the graph-based reasoning initiated by Liu et al [163]. but also extends the concept to the challenges of recognizing multiple objects within an image. Following this thread, Lu et al. [165] first apply the GCN to image semantic segmentation with their Graph-FCN, an approach that redefines the task as a graph node classification problem, successfully integrating local location information often overlooked by standard deep learning methods.
V-E2 Video Analytics
Video analytics with GI models are mainly for HAR and Multi-Object Tracking (MOT). While HAR has been thoroughly explored in Section V-C2, this section delves into the latter, focusing on the innovative approaches that have emerged.
@@ V-E2 Video Analytics
Braso et al. [166] set a message-passing network that not only integrates learning within the MOT paradigm but also crucially extends it to the data association phase (which is traditionally a bottleneck). Simultaneously, Liu et al. [167] introduce the graph similarity model for MOT, which integrates a unique graph representation of object features and relationships, coupled with a graph matching module, to enhance robustness against occlusion and similar appearances in tracking scenarios. Both of them have independently demonstrated the potential of GI models to transform MOT by focusing on different yet complementary aspects of the problem, i.e., global reasoning and robust feature representation.
V-E3 Point Cloud Processing
Different from images in 2D representations, point clouds are organized in 3D structure and provide richer semantics in describing scenes and objects. For instance, many autonomous vehicles use LiDAR to scan the environment and export point cloud data for perception, as in Fig. 9. Yet the 3D data are still accommodated in a graph representation, and thereby GI models are leveraged for efficient processing.
@@ V-E3 Point Cloud Processing
Wang et al. [168] develop EdgeConv, which enhances the representation power of point clouds by recovering topological information. This method uses dynamic graph construction to capture local geometric features, proving its suitability for high-level tasks in computer graphics and vision. Building on this idea of dynamic feature extraction, Shi et al. [162] propose Point-GNN, a GI model for object detection from LiDAR point cloud, which incorporates an auto-registration mechanism to reduce translation variance and a box merging and scoring operator for accurate detection over multiple vertices.
Extending these advancements, Lin et al. [169] develop 3D-GCN to address the challenges of processing and summarizing information in unstructured point clouds for 3D vision applications. Their approach is particularly effective when dealing with data variations such as shift and scale changes, aligning with the principles seen in EdgeConv. Furthermore, Zhang et al. [170] present the linked dynamic GNN, which tackles the challenge of processing sparse, unstructured point clouds by linking hierarchical features from dynamic graphs. This method optimizes the network’s architecture for more effective point cloud classification and segmentation, showing how linking hierarchical features can avoid the vanishing gradient problem and improve overall performance.
Lessons Learned (Sec. V) EGI has catalyzed a broad range of applications beyond traditional edge AI scenarios. These applications typically follow a general working procedure: the system collects data from components of edge networks and formulates them in graphs. The graph data are ingested into a GI model for graph embedding, along with a subsequent model, e.g., FCN and LSTM, for downstream tasks. The key to building a well-performed EGI application is carefully modeling the graph data and properly selecting the GI model.
VI Edge Networks for Graph Intelligence
Given the applications introduced in Sec. V, this section discusses how edge networks serve GI execution in their multi-tier hierarchy. In the following, we survey various computing systems in different architectures according to their reliance on the cloud. Note that in each architecture, we review both model training and model inference systems, and their EGI ratings thus diverge.
VI-A Federated Graph Learning
@@
Edge-cloud synergy extends cloud resources as remote computing assists in optimizing the performance and efficiency of user-centric computing and data processing. Specifically, it enables certain computing tasks reserved at the edge for immediate processing, while employing the cloud for offloading more resource-intensive tasks with long-term storage and complex data analysis. This hybrid computing paradigm combines the differentiated capabilities of edge and cloud computing, thereby admitting flexible service computing for various applications [171].
For edge computation of GI models, typical workloads in the form of edge-cloud synergy are FL of GI models, i.e., Federated Graph Learning (FGL). FGL is the Level-4 EGI system in the rating taxonomy but is distinct from GI-assisted FL (GFL) as discussed in Sec. VII-D. Although they both combine GL and FL, the targeted model to be federatively trained in FGL is exactly GI models but GFL allows the targeted model to be any model not only GI models. Readers particularly interested in the combination of GI and FL may find dedicated reviews [172, 173] for more information. With respect to the segmentation of the data space, FGL can be divided into horizontal ones (graph topology space) and vertical ones (feature data space).
@@
VI-A1 Horizontal Federated Graph Learning
@@
The first setting considers FGL with explicit links across subgraphs on FL clients, as in Fig. 10. The conventional approach in this setting entails training local GI models within individual clients to initially acquire the local representations or node embeddings of the respective graphs. Subsequently, a Federated Learning (FL) algorithm is applied to facilitate the aggregation process, and the FL server, which is usually a cloud server, accumulates the model parameters or gradients received from the clients and performs FL aggregation. The resultant updated parameters are then returned to the clients for subsequent rounds of training. The introduction of FGL within this framework is contingent upon the specific research problems they address. For Non-IID graph data issues, many approaches intend to enhance the adaptation performance of local models or acquire a robust global model through model adjustment like model interpolation [174, 175] and meta learning [176].
The second setting assumes that certain links connecting vertices across different clients are implicit and absent, illustrated in Fig. 10. For the case where the implicit links connect vertices with dissimilar IDs, a prevalent strategy involves rectifying the local graphs by reconstructing the missing links. This ensures the local graph’s completeness, thereby facilitating high-quality graph representation. Furthermore, the inclusion of links connecting clients aids in mitigating the challenges associated with non-IID data. FASTGNN [177] introduces a straightforward link generator in the FL server, which is responsible for reconstructing absent links between clients by introducing Gaussian randomly generated links. These reconstructed edges are broadcast to all clients during FL training, prompting them to update their local graphs. In a similar vein, the link generator in FedGL [68] produces a global pseudo graph with vertex embeddings provided by clients. Subsequently, this global pseudo-graph is disseminated to all clients, enabling them to amend their local graphs to facilitate GI model training. Based on the simple yet effective idea of link generation, follow-up works further develop more complex techniques for better performance [178, 179], consider privacy issues with security means [180, 181, 182], and explore to improve communication efficiency [183, 184].
@@
VI-A2 Vertical Federated Graph Learning
@@
In the first case, graph topology, vertex features, and vertex labels are distributed across different clients. Consequently, clients may not be able to obtain the full graph data with respect to both vertex features and graph topology, as shown in Fig. 11(a). To orchestrate these clients in such semi-blind situations, SGCNN [190] computes a similarity matrix based on the dynamic time-warping algorithm to sketch the graph’s topology without knowing the underlying structure. One-hot encoding is employed to describe vertices’ features under privacy requirements, which will be transmitted to the server to train the global GI model for GI-based tasks. FedSGC [191] considers a two-client decentralized setting with vertices’ features and graph topology being respectively possessed. In this scenario, FedSGC proposes an additively Homomorphic Encryption (HE) method to ensure privacy preservation during FGL.
@@
The second case entails the ownership of solely disparate vertex feature spaces by clients, while the graph topology remains accessible to all clients (Fig. 11(b)). For the case with only two clients and a central server, FedVGCN employs HE to encode intermediate data transferred across clients, where the server generates encryption key pairs and performs global FL aggregation. FMLST [192] assumes a shared global spatio-temporal pattern across clients and applies an MLP to fuse local patterns and global patterns for personalizing client models. Graph-Fraudster [193] investigates adversarial attacks on this vertical FGL scenario and shows differential privacy mechanisms and top-k mechanisms as two viable defense strategies against such attacks.
VI-B Distributed Edge Collaboration
@@
VI-B1 Edge-Cloud Collaboration
@@
Edge-cloud collaboration differs from traditional integrated cloud offloading by splitting computation workload between edge platforms and cloud servers and thus can fortify the duty of the edge side and alleviate the reliance on the remote cloud [16, 198, 199]. This allows edge facilities to be a key processor that blocks sensitive raw data from the cloud and eliminates redundant data transmission in the backhaul. Following this technical path, Branchy-GNN [200] concentrates on GI-based point cloud processing and applies edge-cloud collaboration with a learning-based source-channel coding scheme for bandwidth-efficient intermediate data transfer. It also borrows the idea of joint split learning and early exiting to reduce the computational cost on edge devices [16, 201]. Pyramid proposes a hierarchical architecture for edge-cloud synergetic GNNs that trains a local, small model on edge servers with its collected data and a global model on a cloud server to capture inter-region spatio-temporal relationships. During the runtime, Pyramid [202] allows GI-based predictions with either the local model or the global model in an on-demand fashion and achieves lower inference latency against baselines. Branchy-GNN is a Level-3 EGI system since it only tackles GNN inference at the edge, while Pyramid rates Level-4 because of its joint training and inference designs.
VI-B2 Multi-Edge Collaboration